# Non-Homophilic Graph Pre-Training and Prompt Learning

### v1

@@ Abstract.
Graphs are ubiquitous for modeling complex relationships between objects across various fields. Graph neural networks (GNNs) have become a mainstream technique for graph-based applications, but their performance heavily relies on abundant labeled data. To reduce labeling requirement, pre-training and prompt learning has become a popular alternative. However, most existing prompt methods do not differentiate homophilic and heterophilic characteristics of real-world graphs. In particular, many real-world graphs are non-homophilic, not strictly or uniformly homophilic with mixing homophilic and heterophilic patterns, exhibiting varying non-homophilic characteristics across graphs and nodes. In this paper, we propose ProNoG, a novel pre-training and prompt learning framework for such non-homophilic graphs. First, we analyze existing graph pre-training methods, providing theoretical insights into the choice of pre-training tasks. Second, recognizing that each node exhibits unique non-homophilic characteristics, we propose a conditional network to characterize the node-specific patterns in downstream tasks. Finally, we thoroughly evaluate and analyze ProNoG through extensive experiments on ten public datasets.
@@ 2. Related Work
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018; Xu et al., 2019; Hamilton et al., 2017; Yu et al., 2023) are mainstream technique for graph representation learning. They typically operate on a message-passing framework, where nodes iteratively update their representations by aggregating messages received from their neighboring nodes. However, the effectiveness of GNNs heavily relies on abundant task-specific labeled data and requires re-training for various tasks. Inspired by the success of pre-training methods in the language (Dong et al., 2019; Brown et al., 2020; Gao et al., 2021; Schick and Schütze, 2021) and vision (Bao et al., 2022; Zhou et al., 2022b, a; Zang et al., 2022) domains, pre-training methods (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021; Qiu et al., 2020; Velickovic et al., 2019; You et al., 2020; Tan et al., 2024) have been widely explored for graphs. These methods first pre-train a graph encoder based on self-supervised tasks, then transfer prior knowledge to downstream tasks. However, all these GNNs and pre-training methods are based on the homophilic assumption, overlooking that real-world graphs are generally non-homophilic.
@@ 6. Experiments
Table 1. Summary of datasets.
Wisconsin 1 0.21 - 251 199 1,703 5
Squirrel 1 0.22 - 5,201 217,073 2,089 5
Chameleon 1 0.23 - 2,277 36,101 2,325 5
Cornell 1 0.30 - 183 295 1,703 5
PROTEINS 1,113 0.66 2 39.06 72.82 1 3
ENZYMES 600 0.67 6 32.63 62.14 18 3
Citeseer 1 0.74 - 3,327 4,732 3,703 6
Cora 1 0.81 - 2,708 5,429 1,433 7
BZR 405 - 2 35.75 38.36 3 -
COX2 467 - 2 41.22 43.45 3 -
Table 2. Accuracy evaluation on few-shot node classification.
@@
Table 3. Accuracy evaluation on few-shot graph classification.
@@ 6.2. Few-shot Performance Evaluation
One-shot performance. We present the results of one-shot node and graph classification tasks on non-homophilic graphs in Tables 2 and 3, respectively. We make the following observations: (1) ProNoG surpasses all baseline methods across all settings, outperforming the best competitor by up to 21.49% on node classification and 6.50% on graph classification. These results demonstrate its effectiveness in learning prior knowledge from non-homophilic graphs and capturing nodes’ specific patterns. (2) Other graph prompt learning methods, i.e., GPPT, GraphPrompt, and GraphPrompt+, significantly lag behind ProNoG. Their suboptimal performance can be attributed to their inability to account for a variety of node-specific patterns. These results underscore the importance of our conditional prompting in characterizing node embeddings to capture nodes’ specific patterns. (3) GPPT is at best comparable to, and often performs worse than other baselines because it is not specifically designed for few-shot learning.
@@ C. Further Descriptions of Datasets
We conduct experiments on ten benchmark datasets.

### v4

@@ Abstract.
Graphs are ubiquitous for modeling complex relationships between objects across various fields. Graph neural networks (GNNs) have become a mainstream technique for graph-based applications, but their performance heavily relies on abundant labeled data. To reduce labeling requirement, pre-training and prompt learning has become a popular alternative. However, most existing prompt methods do not differentiate homophilic and heterophilic characteristics of real-world graphs. In particular, many real-world graphs are non-homophilic, not strictly or uniformly homophilic with mixing homophilic and heterophilic patterns, exhibiting varying non-homophilic characteristics across graphs and nodes. In this paper, we propose ProNoG, a novel pre-training and prompt learning framework for such non-homophilic graphs. First, we analyze existing graph pre-training methods, providing theoretical insights into the choice of pre-training tasks. Second, recognizing that each node exhibits unique non-homophilic characteristics, we propose a conditional network to characterize the node-specific patterns in downstream tasks. Finally, we thoroughly evaluate and analyze ProNoG through extensive experiments on ten public datasets. Codes are available at https://github.com/Jaygagaga/ProNoG/tree/main.
@@ 2. Related Work
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018; Fang et al., 2023a; Bo et al., 2023; Fang et al., 2022; Li et al., 2024a, b) are mainstream technique for graph representation learning. They typically operate on a message-passing framework, where nodes iteratively update their representations by aggregating messages received from their neighboring nodes (Xu et al., 2019; Hamilton et al., 2017; Yu et al., 2023; Fang et al., 2023c, b; Jiang et al., 2024a). However, the effectiveness of GNNs heavily relies on abundant task-specific labeled data and requires re-training for various tasks. Inspired by the success of pre-training methods in the language (Dong et al., 2019; Brown et al., 2020; Gao et al., 2021; Schick and Schütze, 2021) and vision (Bao et al., 2022; Zhou et al., 2022b, a; Zang et al., 2022) domains, pre-training methods (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021; Qiu et al., 2020; Velickovic et al., 2019; You et al., 2020; Tan et al., 2024; Mo et al., 2023, 2024; Jiang et al., 2024b) have been widely explored for graphs. These methods first pre-train a graph encoder based on self-supervised tasks, then transfer prior knowledge to downstream tasks. However, all these GNNs and pre-training methods are based on the homophilic assumption, overlooking that real-world graphs are generally non-homophilic.
@@ 6. Experiments
Table 1. Accuracy evaluation on one-shot node classification.
@@
Table 2. Accuracy evaluation on one-shot graph classification.
@@ 6.2. Few-shot Performance Evaluation
One-shot performance. We present the results of one-shot node and graph classification tasks on non-homophilic graphs in Tables 1 and 2, respectively. We make the following observations: (1) ProNoG surpasses all baseline methods across all settings, outperforming the best competitor by up to 21.49% on node classification and 6.50% on graph classification. These results demonstrate its effectiveness in learning prior knowledge from non-homophilic graphs and capturing nodes’ specific patterns. (2) Other graph prompt learning methods, i.e., GPPT, GraphPrompt, and GraphPrompt+, significantly lag behind ProNoG. Their suboptimal performance can be attributed to their inability to account for a variety of node-specific patterns. These results underscore the importance of our conditional prompting in characterizing node embeddings to capture nodes’ specific patterns. (3) GPPT is at best comparable to, and often performs worse than other baselines because it is not specifically designed for few-shot learning.
@@ 7. Conclusions
Acknowledgments
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
@@ B. Homophily and Non-Homophily Methods
Table 6. Summary of datasets.
Wisconsin 1 0.21 - 251 199 1,703 5
Squirrel 1 0.22 - 5,201 217,073 2,089 5
Chameleon 1 0.23 - 2,277 36,101 2,325 5
Cornell 1 0.30 - 183 295 1,703 5
PROTEINS 1,113 0.66 2 39.06 72.82 1 3
ENZYMES 600 0.67 6 32.63 62.14 18 3
Citeseer 1 0.74 - 3,327 4,732 3,703 6
Cora 1 0.81 - 2,708 5,429 1,433 7
BZR 405 - 2 35.75 38.36 3 -
COX2 467 - 2 41.22 43.45 3 -
@@ C. Further Descriptions of Datasets
We conduct experiments on ten benchmark datasets. We summary these datasets in Table 6.
@@ H. Hyperparameter Analysis
I. Further Comparison with Other Baselines
We conduct further one-shot node classification on Wisconsin and compare with other ”pre-training, prompting” baselines. The results are shown in Table 8. ProNoG consistently outperforms these methods, further demonstrating the effectiveness of ProNoG
Table 8. Accuracy evaluation on one-shot node classification.
Methods Wisconsin
Self-pro (Gong et al., 2024) 29.81 $\pm$ 10.33
PSP (Ge et al., 2024) 33.71 $\pm$ 7.16
Prodigy (Huang et al., 2024) 25.34 $\pm$ 0.63
All-in-one (Sun et al., 2023a) 23.59 $\pm$ 0.10
ProNoG 44.72 $\pm$ 11.93
J. Efficiency Analysis
We compare the prompt-tuning time of ProNoG and a competitive baseline GraphPrompt (Liu et al., 2023a) and show the results in Table 9. Empirically, our prompt tuning takes only marginally longer than the strong baseline GraphPrompt, as shown below (in seconds), while significantly outperforming GraphPrompt in terms of performance. Therefore, this trade-off is acceptable.
Table 9. Prompt-tuning times (seconds) on node classification.
Methods Cora
GraphPrompt 1.003
ProNoG 1.513
K. Performance on Large-scale Dataset
We evaluate ProNoG under the same setting as introduced in Section 6.1 on a large-scale dataset Amazon (Zeng et al., 2019), which consists of 1,598,960 nodes, 132,169,734 edges, and 107 classes. The results are shown in Table 10. We observe that ProNoG outperforms competitive baselines, showing the effectiveness of ProNoG on large-scale dataset.
Table 10. Accuracy evaluation on one-shot node classification.
Methods Amazon
GraphPrompt 3.91
ProNoG 4.56